It's not 7. Please help me solve! (:

Thank you in advance!
1 attempts let Check my work Enter your answer in the provided box. For the simple decomposition reaction AB(g) → A(g) + B(g) rate kIABI2 and k-0.70 /mol s. How long will it take for [ABI to reach 1/3 of its initial concentration of 1.50 M?
Answer #1

Given rate,r=K[AB[^2

So order of the reaction is 2

For a second order reaction (1/A) = (1/Ao)+ kt
Where
Ao = initial concentration = 1.50 M
A = concentration left after time t = Ao/3 = 0.5 M
t = time taken = ?
k = rate constant = 0.70 L/mol-s
Plug the values we get t= 1.9~ 2s